OpenStudy (anonymous):

The equation c = 0.44n gives the cost of buying some 44-cent stamps. Which best states the meaning of c? The cost of the stamps in cents The cost of the stamps in dollars The number of stamps on a sheet The number of stamps purchased

OpenStudy (anonymous):

The cost of the stamps in dollars since the 0.44 represents 44 cents

OpenStudy (anonymous):

@DSS thanks so much ! best response :)

OpenStudy (anonymous):

What is an equation in slope-intercept form of the line that passes through (6, 10) and is parallel to the graph OF \[y = 1/3 x-1\]

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

Are the graphs of -5Y = 2 X + 3 and Y=2/5X+4 parallel, perpendicular, or neither?

OpenStudy (anonymous):

Neither